Q1: How to use incremental rate of return analysis given two alternatives and choosing which one is the best?

Q2:

Case 1: Alternative 1 Alternative 2

    Cost 500 700

Annual Cost . 600 800

Annual Benefit 700 . 900

Case 1: Alternative 1 Alternative 2

Cost 700 500

Annual Cost . 800 600

Annual Benefit 900   700

Based on these two cases, how to I apply incremental return analysis to this, and instead of using incremental rate of return analysis, can we use rate of return analysis on both alternatives and determine which alternative is better?
Answer #1

Q1

First we select base alternative. Base alternative is alternative with lowest initial cost

Then we perform incremental IRR analysis

In order to perform incremental IRR Anslysis, we calculate incremental Cash Flow between the alternatives by subtracting cash flow of lower initial cost alternative from higher cost alternative

then we perform incremental IRR analysis using NPW method by equating NPW equation to zero.

if incremental IRR is greater than MARR then we select the higher initial cost option, if it is less than the MARR then select the lower initial cost option.

Q2

Is case 1, base alternative is alternative 1, as it has lower initial cost

Incremental initial cost (2-1) = 700 - 500 = 200

incremental annual cost (2-1) = 800 - 600 = 200

Incremental annual revenue (2-1) = 900 - 700 = 200

We use these value to set up NPW equation and equate it to zero to find incremental IRR

Is case 2, base alternative is alternative 2, as it has lower initial cost

Incremental initial cost (1-2) = 700 - 500 = 200

incremental annual cost (1-2) = 800 - 600 = 200

Incremental annual revenue (1-2) = 900 - 700 = 200

We use these value to set up NPW equation and equate it to zero to find increment

we cannot use rate of return analysis on both alternatives and determine which alternative is better because we need to calculate incremental return on the incremental investment in higher cost alternative to find the best alternative